How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 3 yards long, and 5 pieces, each 2 yards long, are removed?

Answer by rajagopalan(158)

Total Length of Cloth = 24 Yards
First Cut 3x3 yards...= 9 Yards
Balance after First cut=24-9=15 Yards
Second cut 5x2 yards.........=10 Yards
Balance after Second cut=15-10= 5 yards
Answer=5 yards